For a given number of instructions, assume CPI is increased by 20%, and clock cycle time is decreased by 10%. The program execution time ___________.

a. remains the same

b. increases

c. decreases
Answer #1

CPU Time = I * CPI * T.

I = number of instructions in program;

CPI = average cycles per instruction;

T = clock cycle time

T1 = CPI1 * T1

T2 = (1.2)CPI1 *(0.9)T1

T1/T2 = 1/(1.2 * 0.9)

T2 = (1.08) * T1

So The program execution time

b. increases

If you have any doubts comment below and i am happy to help :)